(1) An electric heater carries a current of 13.5 A when operating at a voltage of 1.20 x 10^2 V. What is the resistance of the heater?

(2) The temperature difference between the inside and the outside of a home on a cold winter day is 57 *F. Express this difference on (a) the Celsius scale and (b) the Kelvin scale.

(3) The two mirrors meet at a right angle. The beam of light in the vertical plane P strikes mirror 1 (variables: 40.0* and 1.25 m). (a) Determine the distance (in m) the reflected light beam travels before striking mirror 2. (b) In what direction (with respect to a line normal to mirror 2) does the light beam travel after being reflected from mirror 2?
1
Expert's answer
2020-02-10T09:25:37-0500

1)


"R=\frac{U}{I}=\frac{120}{13.5}=8.89\Omega"

2) a)


"\Delta T=57\frac{5}{9}=31.67\degree C"

b)


"\Delta T=57\frac{5}{9}=31.67\ K"

3)



a)


"d\sin{40}=1.25"

"d=1.94\ m"

b) The direction of light will be opposite to the initial direction.